# Minimal Subgroups of $\operatorname{GL}_{2}(\mathbb{Z}_{S})$

### v2

@@ 1. Introduction
$\rho_{E,p}:\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\operatorname{GL}_{2}(\mathbb{Z}_{p}).$
Recently there has been great interest in studying the groups of the form $G_{E,p}=\operatorname{Im}(\rho_{E,p})$. A few classical results about $G_{E,p}$ include that $\det(G_{E,p})=\mathbb{Z}_{p}^{\times}$ and if $E$ does not have complex multiplication, then $[\operatorname{GL}_{2}(\mathbb{Z}_{p}):G_{E,p}]$ is finite (see Chapter IV of [17]).
Given these two things, a natural question to ask is if there is an elliptic curves $E/\mathbb{Q}$ and prime $p$ such that $G_{E,p}$ has the property that if $H\subsetneq G_{E,p}$, then $\det(H)\subsetneq\mathbb{Z}_{p}^{\times}$? Before answering this question, we phrase it more broadly.
@@
So a more general question would be if there is an elliptic curve $E/\mathbb{Q}$ and a set of prime $S\subseteq\mathbb{Z}$ such that $G_{E,S}\subseteq\prod_{p\in S}G_{E,p}\subseteq\operatorname{GL}_{2}(\mathbb{Z}_{S})$ is minimal and if so, can we classify all of, or maybe almost all of, such curves?
@@ 1.1. First Examples
Suppose that $E/\mathbb{Q}$ is an elliptic curve without complex multiplication that only has bad reduction at 2. From the criterion of Néron-Ogg-Shafarevich we know that the extension $\mathbb{Q}(E[2^{\infty}])/\mathbb{Q}$ is only ramified at 2. From the main theorem of [20] we know that there are no such cubic extensions of $\mathbb{Q}$. This, together with the fact that $\operatorname{GL}_{2}(\mathbb{Z}/2\mathbb{Z})\simeq S_{3}$ would force $E$ to have a point of order 2 defined over $\mathbb{Q}$. Besides being able to conclude that $E$ has to have a rational point of order 2, this allows us to conclude that $G_{E,2}$ is in fact a pro 2-group. Now that we know that $G_{E,2}$ is a pro 2-group, we can use a classical result in group theory that says that any proper maximal subgroup of a pro $p$-group is normal and has index $p$. (Here and elsewhere, we require a maximal subgroup of a profinite group to be closed.)
@@
Searching [9] we find that there are exactly 8 elliptic curves defined over $\mathbb{Q}$ without complex multiplication that only have bad reduction at 2. These elliptic curves have have labels
@@ 1.2. Statement of results and outline
An outline of the paper is as follows. In Section 3 we prove Theorem 1.2 and discuss a generalization to principally polarized abelian varieties. In Section 4 we prove Theorem 1.3 using ideas from $p$-adic Lie theory. In Section 5 we describe a search to find all the genus 0 minimal groups and then compute models for the corresponding modular curves. The models for these curves were computed using [1], the techniques in [13], and guided by the information available in [9]. All the code for for these computations is available at [6]. Lastly, we prove Proposition 1.4 in Section 6.
@@ 1.3. Acknowledgements
The second author is thankful to Amherst College for hospitality during a visit in September 2022.
@@ 2.1. Elliptic Curves
We call $r$ the rank of $E$ over $K$, and $T$ is the torsion subgroup of $E(K)$ and often denoted $E(K)_{\mathrm{tors}}.$ A classical result in the study of elliptic curves is that if we fix an algebraic closure of $K$, denoted $\overline{K}$, then
@@
$\overline{\rho}_{E,n}\colon\operatorname{Gal}(\overline{K}/K)\to\operatorname{GL}_{2}(\mathbb{Z}/n\mathbb{Z}).$
We call $\overline{\rho}_{E,n}$ the mod $n$ representations associated with $E$. Using the mod $n$ representations associated to $E$ and the appropriate inverse limits, we can define the $p$-adic and adelic Galois representations attached to $E$. We denote these representations
$\rho_{E,p}\colon\operatorname{Gal}(\overline{K}/K)\to\operatorname{GL}_{2}(\mathbb{Z}_{p}),\hbox{ and}$
@@
Here we note these representations all depend on which points we pick as a basis for the $n$-torsion of $E(K)$, and changing the basis replaces the image with a conjugate subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/n\mathbb{Z})$. It is not hard to see that if $H\subseteq\operatorname{GL}_{2}(\mathbb{Z}_{S})$ is minimal, then every conjugate of $H$ in $\operatorname{GL}_{2}(\mathbb{Z}_{S})$ is also minimal.
@@ 2.2. Modular Curves
Modular curves are in important tool for studying elliptic curves and their Galois representations. The points on these curves correspond to elliptic curves whose mod $n$ Galois representations are contained inside of a particular subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/n\mathbb{Z})$ up to conjugation. The goal of this subsection is to give the basic definitions and theorems needed for this paper. The interested reader is encouraged to see [10]. For the remainder of the section $n$ will be an integer greater than 1.
@@
called the $j$-map associate to $G$, such that if $E/K$ is an elliptic curve with $j(E)\not\in\{0,1728\}$, then $\operatorname{Im}\rho_{E,n}$ is conjugate to a subgroup of $G$ if and only if there is a $P\in X_{G}(K)$ such that $j(E)=\pi_{G}(P)$.
In the case that $-I\not\in G$, $X_{G}$ is an algebraic stack and its course space is the same as $X_{\tilde{G}}$. In this case the moduli interpretation is different; if $-I\not\in G$, if the image of $\overline{\rho}_{E,n}\subseteq G$, this does not imply the same for the quadratic twist $E_{D}$. In this situation, if $U$ is the complement in $X_{\tilde{G}}$ of the cusps and preimages of $j=0$ and $1728$, there is a universal elliptic curve $\mathcal{E}\to U$ so that $E$ has image contained in $G$ if and only if there is some (not necessarily unique) $t$ so that $E\cong\mathcal{E}_{t}$ over $K$. For more detail see [16, Section 2 and 5].
@@
A careful reading of the moduli interpretation, one notices that an elliptic curve $E/\mathbb{Q}$ having a corresponding point on $X_{G}(\mathbb{Q})$ does not ensure that $\operatorname{Im}\overline{\rho}_{E,n}$ is conjugate to $G$. It only ensures that $\operatorname{Im}\overline{\rho}_{E,n}$ is conjugate to a subgroup of $G$. That said, if $G$ is a minimal group containing all matrices $\equiv I\pmod{n}$, since $\det\circ\overline{\rho}_{E,n}$ is surjective, it must be that $\operatorname{Im}\overline{\rho}_{E,n}=G.$
@@ 3. Minimality is a 2-adic phenomena
Notice that if $H$ is a minimal subgroup of $\operatorname{GL}_{2}(\mathbb{Z}_{S})$, then every maximal open subgroup of $H$ can be obtained as the inverse image of a subgroup of $\mathbb{Z}_{S}^{\times}$ under the map $\det:H\to\mathbb{Z}_{S}^{\times}$. So if $H$ is minimal, then every maximal subgroup of $H$ is normal in $H$. Further, if we let $p\in S$ , $k\in\mathbb{Z}_{+}$ and $\pi_{p^{k}}:H\to\operatorname{GL}_{2}(\mathbb{Z}/p^{k}\mathbb{Z})$ be the standard component-wise reduction map on $H$, the $\pi_{p^{k}}(H)$ also has the property that all of its maximal subgroups are normal.
@@
[8, Chapter 6, Theorem 3 & Corollary 4] Let $G$ be a finite group. The following are equivalent;
@@
Before we prove the main result of this section, we need one more definition
@@
The assumption that $H$ has finite level is not crucial. It is known by work of Nikolov and Segal [12] that every topologically finitely generated profinite group has the property that finite index subgroup is open.
@@
Since $H$ has finite level we know that there is a $k\in\mathbb{Z}_{+}$ such that $p^{k-1}$ divides the level of $H$, while $p^{k}$ does not. From this, it follows that $M\in\rho_{p^{k}}(H)$ if and only if $M\in\rho_{p^{k-1}}(H)$. In particular, $H_{p^{k}}=\rho_{p^{k}}(H)$ contains all matrices $\equiv I\pmod{p^{k-1}}$.
As we noted before, we know that $H_{p^{k}}$ is a finite nilpotent group. From Theorem 3.2 we have that $H_{p^{k}}$ must be the direct product of its $p$-Sylow subgroups. That is, if we let $P={\rm Syl}_{p}(H_{p^{k}})$ and $Q=\prod_{q\neq p}{\rm Syl}_{q}(H_{p^{k}})$, then $H_{p^{k}}\simeq P\times Q$. So given any $B\in H_{p^{k}}$ we can find $X\in Q$ and $Y\in P$ such that $B=XY$ and we also know that $X$ and $Y$ commute with each other. Since $Y\in{\rm Syl}_{p}(H_{p^{k}})$, we know that the order of $Y$ is a power of $p$ and thus $\det(Y)\equiv 1\bmod p.$
@@
Thus for every $B\in H$, $\det(B)$ is a quadratic residue in $\mathbb{Z}/p\mathbb{Z}$. Since $p$ is an odd prime, not every element of $\mathbb{Z}/p\mathbb{Z}$ is a quadratic residue. This contradicts the assumption that $\det(H)=\mathbb{Z}_{S}^{\times}$
@@ 3.1. A diversion into higher dimensional abelian varieties
Elliptic curves are $1$-dimensional abelian varieties, and much of what is true for elliptic curves is also true for abelian varieties mutatis mutandis. Given the aims of this paper, we will not be able to provide all of the background information here, but the interested reader is encouraged to see [11].
@@
$\overline{\rho}_{A,n}:\operatorname{Gal}(\overline{K}/K)\to\operatorname{Aut}(A[n])\simeq\operatorname{GL}_{2g}(\mathbb{Z}/n\mathbb{Z})$
@@
${\rm Mult}\circ\rho_{A,n}:\operatorname{Gal}(\overline{K}/K)\to(\mathbb{Z}/n\mathbb{Z})^{\times}$
@@
With this in hand, we can see that the argument in the proof of Theorem 1.2 generalizes to ${\rm GSp}_{2g}(\mathbb{Z}_{S})$. In particular, if we assume that $A/K$ is an abelian variety of dimension $g$, $p$ is a prime, and the image of the mod $p^{k}$ Galois representation attached to $A$ contains all matrices in ${\rm GSp}_{2g}(\mathbb{Z}/p^{k}\mathbb{Z})$ that are congruent to the identity modulo $p^{k-1}$, then any $X$ in the image of $\rho_{A,p^{k}}$ with order coprime to $p$ must commute with matrices $I+p^{k-1}Z\in{\rm GSp}_{2g}(\mathbb{Z}/p^{k}\mathbb{Z})$. Writing
@@
Any element in the image of $\rho_{A,p^{k}}$ must have the form $XY$ for some $Y$ with order a power of $p$. This implies that ${\rm Mult}(Y)\equiv 1\bmod p$, and so ${\rm Mult}(XY)={\rm Mult}(X){\rm Mult}(Y)\equiv\lambda^{2}\bmod p$. If $p>2$, this contradicts that ${\rm Mult}\circ\rho_{A,p}$ is surjective, and this is a contradiction.

### v4

@@ 1. Introduction
$\rho_{E,p^{\infty}}:\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\operatorname{GL}_{2}(\mathbb{Z}_{p}).$
Recently there has been great interest in studying the groups of the form $G_{E,p}=\operatorname{Im}(\rho_{E,p^{\infty}})$. A few classical results about $G_{E,p}$ include that $\det(G_{E,p})=\mathbb{Z}_{p}^{\times}$ and if $E$ does not have complex multiplication, then $[\operatorname{GL}_{2}(\mathbb{Z}_{p}):G_{E,p}]$ is finite (see Chapter IV of [18]).
Given these two things, a natural question to ask is if there is an elliptic curve $E/\mathbb{Q}$ and prime $p$ such that $G_{E,p}$ has the property that if $H\subsetneq G_{E,p}$, then $\det(H)\subsetneq\mathbb{Z}_{p}^{\times}$? Before answering this question, we phrase it more broadly.
@@
So a more general question would be if there is an elliptic curve $E/\mathbb{Q}$ and a set of primes $S\subseteq\mathbb{Z}$ such that the image of the $S$-adic Galois representation $\operatorname{Im}\rho_{E,S}=G_{E,S}\subseteq\prod_{p\in S}G_{E,p}\subseteq\operatorname{GL}_{2}(\mathbb{Z}_{S})$ is minimal and if so, can we classify all of, or maybe almost all of, such curves?
@@ 1.1. First Examples
Suppose that $E/\mathbb{Q}$ is an elliptic curve without complex multiplication that only has bad reduction at 2. From the criterion of Néron-Ogg-Shafarevich [19] we know that the extension $\mathbb{Q}(E[2^{\infty}])/\mathbb{Q}$ is only ramified at 2. From the main theorem of [22] we know that there are no such cubic extensions of $\mathbb{Q}$. This, together with the fact that $\operatorname{GL}_{2}(\mathbb{Z}/2\mathbb{Z})\simeq S_{3}$ would force $E$ to have a point of order 2 defined over $\mathbb{Q}$. Besides being able to conclude that $E$ has to have a rational point of order 2, this allows us to conclude that $G_{E,2}$ is in fact a pro 2-group. This is because once we know that $E$ has a rational point of order 2, we know that $[\mathbb{Q}(E[2]):\mathbb{Q}]=1$ or $2$ and $[\mathbb{Q}(E[2^{n+1}]):\mathbb{Q}(E[2^{n}])]$ is always a power of 2. Now that we know that $G_{E,2}$ is a pro 2-group, we can use a classical result in group theory that says that any proper maximal subgroup of a pro $p$-group is normal and has index $p$. (Here and elsewhere, we require a maximal subgroup of a profinite group to be closed.)
@@
Searching [10] we find that there are 8 elliptic curves defined over $\mathbb{Q}$ without complex multiplication that only have bad reduction at 2. These elliptic curves have labels
@@
Remark 1.2.
We know that these are in fact all of the elliptic curves over $\mathbb{Q}$ with bad reduction only at 2. This is because any elliptic curve with only bad reduction at two has conductor equal to a power of 2, but there is a bound on the exponent that can appear on 2 from [20, Theorem 10.2]. Thus any elliptic curve over $\mathbb{Q}$ with bad reduction only at 2 has conductor bounded by 256. So the completeness of the data in [10] allows us to conclude that these are all of them.
@@ 1.2. Statement of results and outline
An outline of the paper is as follows. In Section 3 we prove Theorem 1.3 and discuss a generalization to principally polarized abelian varieties. In Section 4 we prove Theorem 1.4 using ideas from $p$-adic Lie theory. In Section 5 we describe a search to find all the genus 0 minimal groups and then compute models for the corresponding modular curves. The models for these curves were computed using [1], the techniques in [14], and guided by the information available in [10]. All the code for these computations is available at [7]. Lastly, we prove Proposition 1.5 in Section 6.
@@ 1.3. Acknowledgements
We would first like to thank the anonymous referee for their thoughtful comments on a previous version of this paper. The second author is thankful to Amherst College for hospitality during a visit in September 2022.
@@ 2.1. Elliptic Curves
We call $r$ the rank of $E$ over $K$, and $T$ the torsion subgroup of $E(K)$, often denoted $E(K)_{\mathrm{tors}}.$ A classical result in the study of elliptic curves is that if we fix an algebraic closure of $K$, denoted $\overline{K}$, then
@@
$\bar{\rho}_{E,n}\colon\operatorname{Gal}(\overline{K}/K)\to\operatorname{GL}_{2}(\mathbb{Z}/n\mathbb{Z}).$
We call $\bar{\rho}_{E,n}$ the mod $n$ representations associated with $E$. Using the mod $n$ representations associated to $E$ and the appropriate inverse limits, we can define the $p$-adic and adelic Galois representations attached to $E$. We denote these representations
$\rho_{E,p^{\infty}}\colon\operatorname{Gal}(\overline{K}/K)\to\operatorname{GL}_{2}(\mathbb{Z}_{p}),\hbox{ and}$
@@
Here we note these representations all depend on which points we pick as a basis for the $n$-torsion of $E(K)$, and changing the basis replaces the image with a conjugate subgroup. It is not hard to see that if $H\subseteq\operatorname{GL}_{2}(\mathbb{Z}_{S})$ is minimal, then every conjugate of $H$ in $\operatorname{GL}_{2}(\mathbb{Z}_{S})$ is also minimal.
@@ 2.2. Modular Curves
Modular curves are an important tool for studying elliptic curves and their Galois representations. The points on these curves correspond to elliptic curves whose mod $n$ Galois representations are contained inside of a particular subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/n\mathbb{Z})$ up to conjugation. The goal of this subsection is to give the basic definitions and theorems needed for this paper. The interested reader is encouraged to see [11]. For the remainder of the section $n$ will be an integer greater than 1.
@@
called the $j$-map associate to $G$, such that if $E/K$ is an elliptic curve with $j(E)\not\in\{0,1728\}$, then $\operatorname{Im}\bar{\rho}_{E,n}$ is conjugate to a subgroup of $G$ if and only if there is a $P\in X_{G}(K)$ such that $j(E)=\pi_{G}(P)$.
In the case that $-I\not\in G$, $X_{G}$ is an algebraic stack and its course space is the same as $X_{\tilde{G}}$, where $\tilde{G}=\left\langle G,-I\right\rangle$. In this case the moduli interpretation is different; if $-I\not\in G$, if the image of $\bar{\rho}_{E,n}\subseteq G$, this does not imply the same for the quadratic twist $E_{D}$. In this situation, if $U$ is the complement in $X_{\tilde{G}}$ of the cusps and preimages of $j=0$ and $1728$, there is a universal elliptic curve $\mathcal{E}\to U$ so that $E$ has image contained in $G$ if and only if there is some (not necessarily unique) $t$ so that $E\cong\mathcal{E}_{t}$ over $K$. For more detail see [17, Section 2 and 5].
@@
A careful reading of the moduli interpretation, one notices that an elliptic curve $E/\mathbb{Q}$ having a corresponding point on $X_{G}(\mathbb{Q})$ does not ensure that $\operatorname{Im}\bar{\rho}_{E,n}$ is conjugate to $G$. It only ensures that $\operatorname{Im}\bar{\rho}_{E,n}$ is conjugate to a subgroup of $G$. That said, if $G$ is a minimal group containing all matrices $\equiv I\pmod{n}$, since $\det\circ\bar{\rho}_{E,n}$ is surjective, it must be that $\operatorname{Im}\bar{\rho}_{E,n}=G.$
@@ 3. Minimality is a 2-adic phenomena
Notice that if $H$ is a minimal subgroup of $\operatorname{GL}_{2}(\mathbb{Z}_{S})$, then every maximal open subgroup of $H$ can be obtained as the inverse image of a subgroup of $\mathbb{Z}_{S}^{\times}$ under the map $\det:H\to\mathbb{Z}_{S}^{\times}$. This is because if $M_{1}$ and $M_{2}$ were maximal subgroups of $H$ such that $\det(M_{1})=\det(M_{2})\subsetneq\mathbb{Z}_{S}^{\times},$ then $\det^{-1}(\det(M_{1}))=\det^{-1}(\det(M_{2}))$ would be a proper subgroup of $H$ that contains both $M_{1}$ and $M_{2}.$ The maximality of $M_{1}$ and $M_{2}$ then forces $M_{1}=M_{2}.$ So if $H$ is minimal, then every maximal subgroup of $H$ is normal in $H$. Further, if we let $p\in S$ , $k\in\mathbb{Z}_{+}$ and $\pi_{p^{k}}:H\to\operatorname{GL}_{2}(\mathbb{Z}/p^{k}\mathbb{Z})$ be the standard component-wise reduction map on $H$, the $\pi_{p^{k}}(H)$ also has the property that all of its maximal subgroups are normal.
@@
[9, Chapter 6, Theorem 3 & Corollary 4] Let $G$ be a finite group. The following are equivalent;
@@
Before we prove the main result of this section, we need one more definition.
@@
The assumption that $H$ has finite level is not crucial. It is known by work of Nikolov and Segal [13] that every topologically finitely generated profinite group has the property that every finite index subgroup is open.
We are now ready to provide a proof of Theorem 1.3.
@@
Since $H$ has finite level we know that there is a $k\in\mathbb{Z}_{+}$ such that $p^{k-1}$ divides the level of $H$, while $p^{k}$ does not. From this, it follows that $M\in\rho_{p^{k}}(H)$ if and only if $M\in\rho_{p^{k-1}}(H)$. In particular, $H_{p^{k}}=\rho_{p^{k}}(H)$ contains all matrices equivalent to the identity mod $p^{k-1}$.
As we noted before, we know that $H_{p^{k}}$ is a finite nilpotent group. From Theorem 3.2 we have that $H_{p^{k}}$ must be the direct product of its $p$-Sylow subgroups. That is, if we let $P={\rm Syl}_{p}(H_{p^{k}})$ and $Q=\prod_{q\neq p}{\rm Syl}_{q}(H_{p^{k}})$, then $H_{p^{k}}\simeq P\times Q$. So given any $B\in H_{p^{k}}$ we can find $X\in Q$ and $Y\in P$ such that $B=XY$. We also know that $X$ and $Y$ commute with each other in $H_{p^{k}}\simeq P\times Q$ since we have $X\in Q$ and $Y\in P$ Since $Y\in{\rm Syl}_{p}(H_{p^{k}})$, we know that the order of $Y$ is a power of $p$ and thus for some $j$ $\det(Y)^{p^{j}}=\det(Y^{p^{j}})=\det(I)\equiv 1\bmod p.$ Fermat’s little theorem says $\det(Y)^{p^{j}}\equiv 1\bmod p$ forces $\det(Y)\equiv 1\bmod p$.
@@
Thus for every $B\in H$, $\det(B)$ is a quadratic residue in $\mathbb{Z}/p\mathbb{Z}$. Since $p$ is an odd prime, not every element of $\mathbb{Z}/p\mathbb{Z}$ is a quadratic residue. This contradicts the assumption that $\det(H)=\mathbb{Z}_{S}^{\times}$. ∎
Remark 3.6.
One way to make sense of what is happening here is that we are requiring our minimal groups to have surjective determinant. Looking at $\mathbb{Z}_{p}^{\times}$, we see that there is structural difference depending on if $p=2$ or if $p$ is odd. When $p=2$, the group $\mathbb{Z}_{2}^{\times}$ is a 2-group, but when $p$ is odd, $\mathbb{Z}_{p}^{\times}$ is not a $p$-group. This is because $(\mathbb{Z}/p\mathbb{Z})^{\times}$ has size $p-1$. This factor of $p-1$ is what prevents $\mathbb{Z}_{p}^{\times}$ from being a $p$-group when $p$ is odd, but there is no problem here when $p=2.$
@@ 3.1. A diversion into higher dimensional abelian varieties
Elliptic curves are $1$-dimensional abelian varieties, and much of what is true for elliptic curves is also true for abelian varieties being careful to adjust the details where necessary. Given the aim of this paper, we will not be able to provide all of the background information here, but the interested reader is encouraged to see [12].
@@
$\bar{\rho}_{A,n}:\operatorname{Gal}(\overline{K}/K)\to\operatorname{Aut}(A[n])\simeq\operatorname{GL}_{2g}(\mathbb{Z}/n\mathbb{Z})$
@@
${\rm Mult}\circ\bar{\rho}_{A,n}:\operatorname{Gal}(\overline{K}/K)\to(\mathbb{Z}/n\mathbb{Z})^{\times}$
@@
With this in hand, we can see that the argument in the proof of Theorem 1.3 generalizes to ${\rm GSp}_{2g}(\mathbb{Z}_{S})$. In particular, if we assume that $A/K$ is an abelian variety of dimension $g$, $p$ is a prime, and the image of the mod $p^{k}$ Galois representation attached to $A$ is nilpotent and contains all matrices in ${\rm GSp}_{2g}(\mathbb{Z}/p^{k}\mathbb{Z})$ that are congruent to the identity modulo $p^{k-1}$, then any $X$ in the image of $\bar{\rho}_{A,p^{k}}$ with order coprime to $p$ must commute with matrices $I+p^{k-1}Z\in{\rm GSp}_{2g}(\mathbb{Z}/p^{k}\mathbb{Z})$. Writing
@@
Any element in the image of $\bar{\rho}_{A,p^{k}}$ must have the form $XY$ for some $Y$ with order a power of $p$. This implies that ${\rm Mult}(Y)\equiv 1\bmod p$, and so ${\rm Mult}(XY)={\rm Mult}(X){\rm Mult}(Y)\equiv\lambda^{2}\bmod p$. If $p>2$, this contradicts that ${\rm Mult}\circ\bar{\rho}_{A,p}$ is surjective, and this is a contradiction.